Solve the system by addition.
x + y = 1
x – y = 5
Hint: You must find both the x and the y values for full credit

Answer by stanbon(75887)

Solve the system by addition.
x + y = 1
x – y = 5
----------
Adding the two equations you get:
2x=6
x=3
--------
Substitute that into one of the equations to solve for y:
3+y=1
y=-2
-----------
Cheers,
Stan H.